OpenStudy (anonymous):

Write the ratio in lowest terms with whole numbers in the numerator and deminator. 24 yrd to 13 1/3 yrd (show steps please and thanks!)

OpenStudy (anonymous):

best start by multiplying by 3

OpenStudy (anonymous):

so at least you can work with whole numbers. \[24\times 3=72\] \[3\times 13\tfrac{1}{3}=40\]

OpenStudy (anonymous):

then you have \[\frac{72}{40}\] which you can reduce to lowest terms. start by dividing top and bottom by 2

OpenStudy (anonymous):

or if you want to cut to the chase, divide top and bottom by 8 and be done in one step

hero (hero):